OpenStudy (anonymous):

Find the derivative arcsin(t^2)

OpenStudy (anonymous):

If I am correct, the formula is u' / sqrt(1-u^2)

OpenStudy (anonymous):

u=t^2 = 2t/sqrt(1-t^4) I that correct?

OpenStudy (nincompoop):

y?

OpenStudy (anonymous):

y?

OpenStudy (nincompoop):

yeah why?

OpenStudy (anonymous):

Why do I need to solve this problem?

OpenStudy (nincompoop):

why that formula

OpenStudy (anonymous):

I found it in my textbook Derivatives of Inverse Trigonometric Functions: d/dx[arcsinu] = u' / sqrt(1-u^2)

OpenStudy (nincompoop):

then why ask us?

OpenStudy (anonymous):

I want to make sure I am getting the correct answer

OpenStudy (nincompoop):

you are

OpenStudy (anonymous):

Thanks :)